# Inverse Reinforcement Learning with Switching Rewards and History Dependency for Characterizing Animal Behaviors

### v2

@@ 3.4 SWIRL Inference Procedure
It is worth noting that this EM derivation seeks to maximize the log-likelihood of policy $\pi_{z_{t}}(a_{t}|s_{t}^{L};r_{z})$ given the posterior of $z_{t}$ (Eq. 3). This procedure can be incorporated into the Maximum Entropy IRL (MaxEnt IRL) (Ziebart et al., 2008, 2010) framework, since optimizing the reward by maximizing $\log\pi$ corresponds to the Lagrangian dual of the standard MaxEnt IRL approach (Zeng et al., 2022). As a result, SWIRL can leverage a variety of IRL methods within the MaxEnt IRL framework: SWIRL can be integrated with model-based methods to infer the policy used to optimize the reward, such as Soft-Q iteration (Haarnoja et al., 2017), or model-free methods, such as Soft-Q learning (Haarnoja et al., 2017) and Soft Actor-Critic (Haarnoja et al., 2018). Alternatively, we can also convert the MaxEnt IRL problem to a model-free inverse Q-learning problem, such as IQ-Learn (Garg et al., 2021). Although model-free methods are often more general and scalable, we find that the model-based Soft-Q iteration approach is more suitable for our behavioral neuroscience study setting, where environment sizes tend to be small, data availability is limited, and the accuracy requirement for reward recovery is high. Accordingly, we use the Soft-Q iteration approach in our main experiments. Nevertheless, in Appendix D.1 we also demonstrate that SWIRL can be integrated with the model-free IQ-Learn to achieve solid performance, illustrating SWIRL’s potential in broader applications.
@@ 4.2 Application to long, non-stereotyped mouse trajectories
Figure 3: Water-restricted labyrinth experiment. (A) Setup for the labyrinth experiment. (B) Boxplot showing held-out test LL over 10 runs, with the x-axis representing the five different models. Outlier selection method is described in Appendix B.8. (C) History dependency inferred by SWIRL (S-2), as reflected in the reward map for the water mode. (D) Inferred reward maps from SWIRL (S-2) under three hidden modes: water, home, and explore. To enhance visualization, the inferred reward $r(s_{t},s_{t-1})$ was averaged over $s_{t-1}$ to produce $r(s_{t})$. (E) Segments of a trajectory from held-out test data, predicted by four SWIRL models. The orange dot indicates when the mouse visits the water port, while the red cross denotes the mouse’s visit to state 0 (home) at that time. (F) Trajectories segmented into hidden modes based on SWIRL (S-2) predictions.
We then applied SWIRL to the long, non-stereotyped trajectories of mice navigating a 127-node labyrinth environment with water restrictions (Rosenberg et al., 2021). In this experiment, a cohort of 10 water-deprived mice moved freely in the dark for 7 hours. A water reward was provided at an end node (Fig. 3A), but only once every 90 seconds at most. Similar to the simulated experiment, the 90-second condition forces the mice to leave the port after drinking water, leading to a non-Markovian internal reward function. For our analysis, we segmented the raw node visit data into 238 trajectories, each comprising 500 time points. This data format presents a considerably greater challenge compared to the same dataset processed with more handcrafted methods in previous IRL applications (Ashwood et al., 2022a; Zhu et al., 2024), which were limited to clustered, stereotyped trajectories of only 20 time points in length.
@@ 4.2.1 SWIRL inferred interpretable history-dependent reward maps
We applied SWIRL to 80% of the 238 mouse trajectories from the water-restricted labyrinth experiments. According to Rosenberg et al. (2021), mice quickly learned the labyrinth environment and began executing optimal paths from the entrance to the water port within the first hour of the experiment. Therefore, we assume the mice acted optimally concerning the internal reward function guiding their behavior. We tested a number of hidden modes $Z$ from 2 to 5 in Appendix B.6.1 and found $Z=3$ to be the best fit. Fig. 3B displays the held-out test LL for MaxEnt and the SWIRL variations based on the remaining 20% of trajectories. The state dependency in hidden-mode switching dynamics and the history dependency in the reward function contributed to improved test performance. The final SWIRL model (S-2) successfully recovered a water reward map, a home reward map, and an explore reward map (Fig. 3D). For better visualization, we averaged the S-2-recovered history-dependent rewards across previous states and normalized the reward values to a range of (0, 1). In the water reward map, mice received a high reward for visiting the water port. In the home reward map, there was a high reward for visiting state 0 at the center of the labyrinth, which also served as the entrance and exit. Mice occasionally went to state 0 to enter or leave the labyrinth and sometimes passed by on their way to other nodes. In the explore reward map, mice received a high reward for exploring areas of the labyrinth other than state 0 and the water port.
Importantly, SWIRL recovered an interpretable history-dependent reward map for the water port (Fig. 3C). It indicates that mice receive a high reward (1.0) for reaching the water port when their previous location was not the water port. If their prior location was the water port, there is still a reward (0.7) for staying there, but the reward for leaving the water port is even higher (0.9). This observation aligns with the water port design, as mice can only obtain water once every 90 seconds. Consequently, it makes sense that the mice would want to leave the water port after reaching it. Such insights would not be captured by a Markovian reward function that depends solely on the current state.
@@ 4.2.2 SWIRL inferred interpretable history-dependent hidden-mode segments
We then visualized all mouse trajectories based on the hidden-mode segments predicted by SWIRL (S-2) (Fig. 3F). In segments classified as water mode, mice start from various locations in the labyrinth and move toward the water port. In segments identified as home mode, mice begin from distant nodes and head toward the center of the labyrinth (home). In segments categorized as explore mode, mice start from junction nodes or the water port and explore end nodes other than the water port. This result demonstrates that SWIRL can identify sub-trajectories of varying lengths from raw data spanning 500 time points, allowing us to visualize them together and reveal clustered behavioral strategies. This capability has not been achieved by previous studies on freely moving animal behavior over extended recording periods, and we conducted this analysis without prior knowledge of the locations of the water port or home.
We also provide a detailed visualization of the hidden-mode segments from an example trajectory in the held-out test data and compare the segmentation performance of the four SWIRL variations (Fig. 3E). In the S-2 segments, visits to the water port (indicated by orange dots) consistently occur at the end of a water mode segment, while visits to state 0 (home) (indicated by red crosses) typically happen at the conclusion of a home mode segment. Notably, home mode segments that do not include a visit to state 0 can still be valid, as these segments may end at state 1 or 2 (see Appendix C.1). In contrast, the I-1, I-2, and S-1 segments exhibit instances of water segments that do not involve a visit to the water port, along with many home segments that lack clear interpretability. Overall, S-2 successfully identifies robust segments of reasonable length, avoiding the numerous rapid switches seen in the other variations. We attribute this to both the state dependency of hidden mode transitions and the history dependency in rewards. This suggests that mice are unlikely to make quick changes in their decisions; instead, they make choices based on their current location and take into account at least two locations while navigating the maze.
@@ 5 Discussion
Beyond its relevance to neuroscience study, SWIRL also benefits the broader machine learning community. It offers a flexible framework with history-dependent switching reward functions that can be integrated with various model-free IRL algorithms. For example, in Appendix D.1, we highlight SWIRL’s potential to be integrated with IQ-Learn, a state-of-the-art scalable model-free IRL method that has been employed in training large language models (Wulfmeier et al., 2024). Future development of SWIRL that emphasizes scalability, such as incorporating RNNs or transformer-based policy model will be particularly valuable for areas such as robotics and large language models, where adaptable IRL strategies are increasingly in demand.
@@ Appendix C Appendix C
C.1 An example labyrinth trajectory
Figure 8: Hidden mode segments in a labyrinth trajectory. (A) Segments of a trajectory from held-out test data, predicted by SWIRL (S-2). The orange dot indicates when the mouse visits the water port, while the red cross denotes the mouse’s visit to state 0 (home) at that time. (B) The segments of the trajectory shown in (A) are plotted within the labyrinth.
To further explore the hidden mode segments of the trajectory from held-out test data presented in Fig. 3E, we visualized the segments corresponding to each hidden mode in this trajectory in detail (Fig. 8B). The visualization reveals that ”home” segments can remain valid even without a visit to state 0, as these segments often instead terminate at state 1 or state 2, which are next to state 0.
Appendix D Appendix D
D.1 Model-free SWIRL on a simulated gridworld environment

### v3

@@ 3.4 SWIRL Inference Procedure
It is worth noting that this EM derivation seeks to maximize the log-likelihood of policy $\pi_{z_{t}}(a_{t}|s_{t}^{L};r_{z})$ given the posterior of $z_{t}$ (Eq. 3). This procedure can be incorporated into the Maximum Entropy IRL (MaxEnt IRL) (Ziebart et al., 2008, 2010) framework, since optimizing the reward by maximizing $\log\pi$ corresponds to the Lagrangian dual of the standard MaxEnt IRL approach (Zeng et al., 2022). As a result, SWIRL can leverage a variety of IRL methods within the MaxEnt IRL framework: SWIRL can be integrated with model-based methods to infer the policy used to optimize the reward, such as Soft-Q iteration (Haarnoja et al., 2017), or model-free methods, such as Soft-Q learning (Haarnoja et al., 2017) and Soft Actor-Critic (Haarnoja et al., 2018). Alternatively, we can also convert the MaxEnt IRL problem to a model-free inverse Q-learning problem, such as IQ-Learn (Garg et al., 2021). Although model-free methods are often more general and scalable, we find that the model-based Soft-Q iteration approach is more suitable for our behavioral neuroscience study setting, where environment sizes tend to be small, data availability is limited, and the accuracy requirement for reward recovery is high. Accordingly, we use the Soft-Q iteration approach in our main experiments. Nevertheless, in Appendix C.1 we also demonstrate that SWIRL can be integrated with the model-free IQ-Learn to achieve solid performance, illustrating SWIRL’s potential in broader applications.
@@ 4.2 Application to long, non-stereotyped mouse trajectories
Figure 3: Water-restricted labyrinth experiment. (A) Setup for the labyrinth experiment. (B) Boxplot showing held-out test LL over 10 runs, with the x-axis representing the five different models. Outlier selection method is described in Appendix B.8. (C) History dependency inferred by SWIRL (S-2), as reflected in reward maps of the water mode and the explore mode. (D) Inferred reward maps from SWIRL (S-2) under three hidden modes: water, home, and explore. To enhance visualization, the inferred reward $r(s_{t},s_{t-1})$ was averaged over $s_{t-1}$ to produce $r(s_{t})$. (E) Segments of a trajectory from held-out test data, predicted by four SWIRL models. The orange dot indicates when the mouse visits the water port, while the red cross denotes the mouse’s visit to state 0 (home) at that time. (F) Trajectories segmented into hidden modes based on SWIRL (S-2) predictions.
We then applied SWIRL to the long, non-stereotyped trajectories of mice navigating a 127-node labyrinth environment with water restrictions (Rosenberg et al., 2021). In this experiment, a cohort of 10 water-deprived mice moved freely in the dark for 7 hours. A water reward was provided at an end node (Fig. 3A), but only once every 90 seconds at most. Similar to the simulated experiment, the 90-second condition forces the mice to leave the port after drinking water, leading to a non-Markovian internal reward function. For our analysis, we segmented the raw node visit data into 238 trajectories, each comprising 500 time points. This data format presents a considerably greater challenge compared to the same dataset processed with more handcrafted methods in previous IRL applications (Ashwood et al., 2022a; Zhu et al., 2024), which were limited to clustered, stereotyped trajectories of only about 20 time points in length.
@@ 4.2.1 SWIRL inferred interpretable history-dependent reward maps
We applied SWIRL to 80% of the 238 mouse trajectories from the water-restricted labyrinth experiments. According to Rosenberg et al. (2021), mice quickly learned the labyrinth environment and began executing optimal paths from the entrance to the water port within the first hour of the experiment. Therefore, we assume the mice acted optimally concerning the internal reward function guiding their behavior. We tested a number of hidden modes $Z$ from 2 to 5 in Appendix B.6.1 and found $Z=3$ to be the best fit. Fig. 3B displays the held-out test LL for MaxEnt and the SWIRL variations based on the remaining 20% of trajectories. The state dependency in hidden-mode switching dynamics and the history dependency in the reward function contributed to improved test performance. The final SWIRL model (S-2) successfully recovered a water reward map, a home reward map, and an explore reward map (Fig. 3D). For better visualization, we averaged the S-2-recovered history-dependent rewards across previous states and normalized the reward values to a range of (0, 1). In the water reward map, mice received a high reward for visiting the water port. In the home reward map, there was a high reward for visiting state 0 at the center of the labyrinth, which also served as the entrance and exit. Mice occasionally went to state 0 to enter or leave the labyrinth and sometimes passed by on their way to other nodes. In the explore reward map, mice received a high reward for exploring areas of the labyrinth other than state 0.
Importantly, SWIRL recovered an interpretable history-dependent reward map for the water port (Fig. 3C). In the water mode, mice always receive a high reward for reaching the port, irrespective of where they were before. On the contrary, in the explore mode the reward depends on their previous location: if the mouse was already at the water port, leaving it is valued more (1.0) than staying (0.4); if the mouse came from elsewhere, reaching the port still provides a substantial reward (0.7). This observation aligns with the water port design, as mice can only obtain water once every 90 seconds. Consequently, it makes sense that the mice would want to leave the water port after reaching it.
@@ 4.2.2 SWIRL inferred interpretable history-dependent hidden-mode segments
We then visualized all mouse trajectories based on the hidden-mode segments predicted by SWIRL (S-2) (Fig. 3F). In segments classified as water mode, mice start from various locations in the labyrinth and move toward the water port. In segments identified as home mode, mice begin from distant nodes and head toward the center of the labyrinth (home). In segments categorized as explore mode, mice start from junction nodes or the water port and explore end nodes. This result demonstrates that SWIRL can identify sub-trajectories of varying lengths from raw data spanning 500 time points, allowing us to visualize them together and reveal clustered behavioral strategies. This capability has not been achieved by previous studies on freely moving animal behavior over extended recording periods, and we conducted this analysis without prior knowledge of the locations of the water port or home.
We also provide a detailed visualization of the hidden-mode segments from an example trajectory in the held-out test data and compare the segmentation performance of the four SWIRL variations (Fig. 3E). In the S-2 segments, visits to the water port (indicated by orange dots) consistently occur near the end of a water mode segment, while visits to state 0 (home) (indicated by red crosses) typically happen at the conclusion of a home mode segment. Notably, home mode segments that do not include a visit to state 0 can still be valid, as these segments may end at state 1 or 2. In contrast, the I-1, I-2, and S-1 segments exhibit instances of water segments that do not involve a visit to the water port, along with many home segments that lack clear interpretability. Overall, S-2 successfully identifies robust segments of reasonable length, avoiding the numerous rapid switches seen in the other variations. We attribute this to both the state dependency of hidden mode transitions and the history dependency in rewards. This suggests that mice are unlikely to make quick changes in their decisions; instead, they make choices based on their current location and take into account at least two locations while navigating the maze.
@@ 5 Discussion
Beyond its relevance to neuroscience study, SWIRL also benefits the broader machine learning community. It offers a flexible framework with history-dependent switching reward functions that can be integrated with various model-free IRL algorithms. For example, in Appendix C.1, we highlight SWIRL’s potential to be integrated with IQ-Learn, a state-of-the-art scalable model-free IRL method that has been employed in training large language models (Wulfmeier et al., 2024). Future development of SWIRL that emphasizes scalability, such as incorporating RNNs or transformer-based policy model will be particularly valuable for areas such as robotics and large language models, where adaptable IRL strategies are increasingly in demand.
@@ Appendix C Appendix C
C.1 Model-free SWIRL on a simulated gridworld environment